x2 - 12x + 32 = 0, I cant figure it out. please help me
Found 2 solutions by ewatrrr, edjones:
Answer by ewatrrr(24785)

 
Hi,
x^2 - 12x + 32 = 0 Factor
(x-8)(x-4) = 0, x = 4, 8
Check Factoring using FOIL
F First terms
O Outside terms -4x
I Inside terms -8x
L Last terms

Answer by edjones(8007)

x2 - 12x + 32 = 0
(x-8)(x-4)=0 Factor
x-8=0
x=8,
x-4=0
x=4
.
Ed